OpenStudy (anonymous):

Can you help me for this? A tree needs to be staked down before a storm. If the ropes can be tied on the tree trunk 17 feet above the ground and the staked rope should make a 60° angle with the ground, how far from the base of the tree should each rope be staked? Round to the nearest foot?

OpenStudy (anonymous):

17/base = tan 60 base = 17/ tan 60

OpenStudy (anonymous):

17/base = tan 60 base = 17/ tan 60

OpenStudy (anonymous):

17/base = tan 60 base = 17/ tan 60

OpenStudy (anonymous):

\[17 \cot (60 {}^{\circ})=\frac{17}{\sqrt{3}}=9.81495 \]